The asteroid 243 Ida has a mass of about 4.0×1016kg and an average radius of about 16 km (it’s not spherical, but you can assume it is).

Calculate the speed with which you would have to throw a rock to put it into orbit around the asteroid 243 Ida, near the surface.

How long would it take for your rock to return and hit you in the back of the head?
Answer #1

The orbit radius and asteroid radius would be approximately the same = R
Centripetal force (on rock mass m) is provided by grav attraction ..

mv²/R = G.Ma.m / R²
v² = G.Ma /R .. .. 6.67^-11 x 4.0^16kg / 16.0^3m = 167 ..

v = √167 .. .. ►v = 12.90 m/s .. (13.0 m/s)

b) Time = Distance/velocity = 2pi*r/v = (2pi*16000)/13 = 7733.15 sec